As everyone knows, bullets bounce from Superman’s chest. Suppose Superman, mass
104 kg, while not moving, is struck by a 4.2 g bullet moving with a speed of 835 m/s. If the
collision is elastic, find the speed that Superman had after the collision. (Assume the
bottoms of his superfeet are frictionless.)
1
Expert's answer
2011-08-25T13:26:15-0400
Let m1 be the mass of the bullet (m1 = 4.2 g), m2 - the mass of Superman (m2 = 104 kg), .U - speed of the bullet before the colision,
V - speed of Superman after the collision.
Consider that the bullet stopped after the collision.
According to the law of conservation of linear momentum:

m1U = m2VV = U*m1/m2V = 835*0.0042/104 = 0.0337 m/s

So, the speed of Superman is 0.0337 m/s.